Calculate and express your answer to the correct number of significant figures

(1.00866 - 1.00728)/6.02205 x 10^23
Answer #1

During addition and subtraction, the answer cannot have more digits to the right of the decimal point than either of the original numbers.

Therefore 1.00866-1.00728 = 0.00138 (the given two numbers have same number number of decimal places).

During multiplication and division, the number of significant figures in the final product or quotient is determined by the original number that has the smallest number of significant figures.

Now, 0.00138/6.02205*10^23 = 0.000229 * 10^-23 = 2.29 * 10^-27(0.00138 has 3 significant figures and 6.02205 has six significant figures)

all 0.00138, 0.000229 and 2.29 have same number of significant figures.